# Josephson effect in a Fibonacci quasicrystal

### v1

@@ V.2 Topological invariant from critical current
As shown in Fig. 14(c), the subgap states in the Fibonacci chain wind across the gap with the phason angle $\phi$ [17] (see also discussion in Sec. II.2). Our idea to measure the winding is based on that the closer these states wind to the Fermi level, the stronger their contribution is to the current-phase relation, and thus the critical current. Specifically, by tuning the chemical potential close to a gap with winding number $|q|$, the variation with $\phi$ in the spectrum comes predominantly from the winding state, since all other states in comparison show no significant winding (per definition) and therefore have a constant contribution as function of $\phi$. Furthermore, the winding state is at its closest and furthest from the Fermi level exactly $|q|$ times. The conjecture is therefore that this should in turn produce an oscillation with period $|q|$ in the current. We note that this should furthermore even be observable in a small region outside the gap.
@@
To demonstrate the measurement the winding numbers $|q|$, we plot in Fig. 15(a) the critical current ${I_{\mathrm{c}}}$ as a function of $\phi$ at several $\mu$ (line colors) inside the topological gaps, illustrating that ${I_{\mathrm{c}}}$ has clearly distinct number of oscillation periods $p$ with $\phi$. In contrast, Fig. 15(b) illustrates that between the gaps there is no distinct oscillation period [see Fig. 14(c) for gap locations]. To more clearly highlight the oscillation periods $p$, we compute the power spectrum $P$ via the discrete Fourier transform $\mathcal{F}$ with respect to $\phi$
@@
which we normalize with $\mathcal{N}=\sum_{p}P(\mu)$. We also subtract the mean critical current to avoid the trivial peak at $p=0$ in the power spectrum, but refrain from further signal processing operations (e.g. windowing). Figure 15(c) shows the power spectrum when $\mu$ is inside the gaps (solid), illustrating a single well-defined peak at the period corresponding to the expected winding number, $p=|q|$ (dashed vertical lines). Barely visible are additional peaks at integer multiples of the main peak as well as at other periods, e.g. spurious peaks due to the discrete Fourier transform. In contrast, Fig. 15(d) shows that there is no single well-defined peak in the power spectrum when $\mu$ is outside the gaps, but instead a number of small peaks of roughly equal size. Coincidentally, we note that these peaks occur at the Fibonacci numbers $F_{n}\in\{1,2,3,5,8,13,\ldots\}$, with additional smaller peaks at integer multiples of these “Fibonacci peaks” due to the discrete Fourier transform. We consistently find these Fibonacci peaks at other values of $\mu$ and for other junctions (see further below), but have no clear explanation for their appearance, other than that the Fibonacci numbers and golden ratio have a tendency to appear throughout various quantities in the Fibonacci chain [71, 17]. Next, we note that when the Fermi level lies in the middle of the largest gaps ($|q|=1,2,3$), ${I_{\mathrm{c}}}$ vanishes when the state lies at the gap edge, while ${I_{\mathrm{c}}}$ is finite when the winding state is inside the gap, see for instance $\mu=1.57{t_{\mathrm{B}}}$ for the $|q|=3$ gap in Fig. 15(a). These scenarios correspond to when either no state or only the winding state is close to the Fermi level, respectively, thus demonstrating that the winding states can fully carry the critical current.
Figure 15: (a) Critical current ${I_{\mathrm{c}}}$ as a function of phason angle $\phi$ in a $C_{9}$ junction (no repetition) with ${t_{\mathrm{int}}}={t_{\mathrm{B}}}$, $\rho=0.8$, ${\rho_{\mathrm{S}}}=1$, when the chemical potential $\mu$ lies inside a gap of the Fibonacci chain, and (b) outside the gaps. (c) Power spectrum $P$ of ${I_{\mathrm{c}}}$ (solid) of (a) with additional $\mu$, with a clear period $p$ corresponding to the winding number $|q|$ of each gap, $p=|q|$ (dashed). (d) Power spectrum of (b) and Fibonacci numbers $F_{n}$ (dashed). $\mathcal{N}$ is a normalization constant.
Figure 16: (a) Power spectrum $P$ of the critical current ${I_{\mathrm{c}}}$ as a function of chemical potential $\mu$ and oscillation period $p$ in the $C_{9}$ junction (no repetition) at $\rho=0.8$, ${\rho_{\mathrm{S}}}=1$, ${t_{\mathrm{int}}}=0.5{t_{\mathrm{B}}}$, with normalization $\mathcal{N}$. (b) Maximum of the power spectrum at each $\mu$. Cyan lines: winding numbers $|q|=p$, with length corresponding to each gap width.
The above discussion of Fig. 15 showed just few select examples to illustrate the winding states cause a distinct oscillation with period $p=|q|$ in the critical current. Next we show in Fig. 16(a) the corresponding full power spectrum between half-filling and the band edge, illustrating the same three kinds of peaks, namely the main peaks from winding numbers $|q|$ when $\mu$ is inside the gaps (cyan lines), the Fibonacci peaks outside the gaps, and additional spurious peaks due to the discrete Fourier transform. This is more clearly illustrated in Fig. 16(b) where we plot $\mathrm{max}\{P,p\}$, i.e. only the largest peak at each $\mu$. All gaps with winding numbers $|q|<20$ are clearly distinguishable. Furthermore, the signature of the largest gaps extends slightly beyond the gap itself. This is one reason why the $|q|=20$ gap is not visible in Fig. 16(b) since it lies too close to the largest gap $|q|=1$, although it is clearly visible in Fig. 16(a) as the largest peak for $p=20$. Apart from the Fibonacci peaks appearing outside the gaps, the only additional peak found is the period doubling, e.g. when the Fermi level lies inside the gaps with $|q|\leq 4$. We expect that a significantly better result can be obtained by applying refined signal processing operations, and that peak fitting can yield a better quantification of the uncertainty and spurious peaks. Still, our results show that even our simple approach can be used to infer the topological invariant and winding number from the critical current.
Finally, for the sake of completeness, we comment on the above procedure for other parameter values and Fibonacci chains. We generally find that close to ${t_{\mathrm{int}}}={t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$, the spurious peaks due to the discrete Fourier transform (e.g. period doubling) become smaller, while the Fibonacci peaks are quite pronounced. Further from ${t_{\mathrm{int}}}={t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$, the situation is reversed, i.e. with larger influence of the spurious peaks, and smaller influence of the Fibonacci peaks. Importantly, the peaks due to the winding numbers $p=|q|$ of the topological gaps are clearly distinguishable in both cases, where in addition these peaks stretch further outside the gaps further from ${t_{\mathrm{int}}}={t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$. Additionally, we find that by comparing the results for different parameter values, we can more reliably capture the winding numbers in the smallest gaps. We find a similar conclusion also for $\rho=0.5$, $\rho=1.2$ and $\rho=1.5$. We also find that the procedure works with and without repetition of the Fibonacci approximant, as well as in longer Fibonacci approximants since these show the same major gap structure and winding numbers. The instances when the procedure works less well are e.g. in too short approximants due less clear winding with $\phi$, or when the hopping ratio becomes either very small or very large, or very far from ${t_{\mathrm{int}}}={t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$, e.g. as ${t_{\mathrm{int}}}\to 0$, since the spurious peaks become so large that it can be difficult to reliably identify the winding numbers.
@@ VI Concluding remarks
We study the influence of quasiperiodicity on the DC Josephson effect, by considering a ballistic hybrid superconductor-quasicrystal-superconductor Josephson junction. We consider the Fibonacci chain as a quasicrystal model system, from the short junction limit with just a few atomic sites, to the long junction limit with several thousand sites. Furthermore, we study Fibonacci chains both without and with repetition, thus essentially modelling both quasicrystals and their approximants [21], embedded across two superconducting leads. As potential experimental realizations of these models, we propose a quasiperiodic engineered atomic chain [115] or metamaterial [155, 156], or moiré structure [118, 119, 120, 121], embedded across two bulk superconductors. Alternatively, we propose a 3D generalization with superconductors sandwiching a layered material, the latter consisting of 2D periodic lattices stacked according to the Fibonacci sequence along the third dimension [122, 123, 124, 125, 126] along which a Josephson current is applied. It was recently shown that superconductivity in the 1D Fibonaci chain extend to such a 3D scenario [70].
@@
We study the Josephson effect via the low-energy ABS spectrum, current-phase relation, and critical current. We exhaustively investigate how these depend on the superconducting phase difference, quasiperiodic degrees of freedom, hopping parameters, chemical potential, and junction length. We find that although the current-phase relation is still $2\pi$ periodic, with either a sinusoidal or sawtooth profile, quasiperiodicity leads to ABS with quasiperiodic oscillations in their probability density, also at perfect resonance. Importantly, we find that quasiperiodicity qualitatively modifies the condition for emergent zero-energy ABS and maximal critical current. We demonstrate that this condition changes between a few simple functional forms in a quasiperiodic manner as the junction length increases. Consequently, we find that the critical current shows quasiperiodic oscillations as a function of junction length, on top of the monotonic decay also found in crystalline junctions [142]. Based on these results we conclude that there might be large sample-to-sample fluctuations between different quasiperiodic junctions, depending on the microscopic details.
@@
Surprisingly, despite proposals for enhanced superconductivity and proximity effect [66, 67, 68, 69, 70, 73, 71, 72], we find that the critical current is not generally enhanced, especially compared to a crystalline junction in the perfect ballistic limit and with a zero-energy state. However, beyond this idealized scenario, we find that quasiperiodicity can significantly enhance the critical current, especially at reduced coupling to the superconducting leads.
@@
We find that by varying the chemical potential in the junction, the junction changes between between SNS and SIS behavior in a fractal manner, due to the intrinsic fractal energy spectrum and topological gap structure of the Fibonacci chain [17]. Each topological gap hosts a subgap state that winds as a function of the phason angle $\phi$, both across the gap and in real space across the Fibonacci chain, with winding number $|q|$ given by a gap labeling theorem [43, 44, 45, 46]. We find that such winding of the subgap states leads to the critical current oscillating with the winding number when the chemical potential is tuned through the gaps, while outside the gaps the critical current oscillates according to the Fibonacci numbers. We therefore demonstrate how the critical current can in principle measure topological invariant in the Fibonacci chain. In summary, these results show how Josephson junctions can be used to probe the intricate physics of quasiperiodic systems, including their interplay with ordered states such as superconductivity.
@@
As an outlook, there are many open questions such as the influence of quasiperiodicity on the AC Josephson effect, different heterostructures and sample realizations, the influence of defects and disorder, finite temperature, as well as different strengths or symmetries of the superconducting order parameter. It would also be interesting to find a connection with the topological invariant without relying on the phason angle [59]. Furthermore, other theoretical models like the scattering matrix approach, or analytic calculations using perturbation theory and renormalization group theory [59], might shine additional light on the rich physics of these systems. Beyond quasicrystals, fractal lattices are another example of interesting aperiodic structures to study the interplay of proximitized phase-coherent phenomena [7, 8, 9]. We further note that the supercurrent-magnetic field relation was recently studied in the transverse direction through a single layer of such a fractal [157], i.e. no current was transmitted along fractal degree of freedom itself. Studying the Josephson effect and currents propagating along a fractal would thus also be an interesting outlook.
@@ VII Acknowledgements
Appendix A Condition for zero-energy ABS
@@
In this Appendix, we demonstrate how we obtain the condition ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ for the emergence of zero-energy ABS, first in the Fibonacci approximants $C_{n}$ with lengths $L=a_{0}F_{n}$ presented in Table 1, then for more general Fibonacci chains of arbitrary discrete lengths $L$ modeled by the characteristic function Eq. (2), thus supplementing the discussion in Sec. IV.1 and Sec. IV.2. Specifically, we vary simultaneously the hopping parameters ${t_{\mathrm{int}}}$, $\rho\equiv{t_{\mathrm{A}}}/{t_{\mathrm{B}}}$ and ${\rho_{\mathrm{S}}}\equiv{t_{\mathrm{S}}}/{t_{\mathrm{B}}}$ and look for when the lowest-energy ABS reaches zero energy non-trivially (i.e. ignoring trivial cases such as $\rho\to 0$ or ${t_{\mathrm{int}}}\to 0$). We parametrize the phase space with zero-energy ABS as ${t_{\mathrm{int}}}={t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$, where we find a sawtooth current-phase relation and the largest critical current. In a Fibonacci chain with even number of sites, the zero-energy ABS at ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ occurs due to perfect Andreev reflection (zero normal reflection) [141], also with a perfect wave function matching at the interface as evidenced by the probability density of the wave function. In a Fibonacci chain with odd number of sites, there is already a zero-energy state in the normal state [151] (see Sec. III.2), and ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ corresponds to the parameter space where this state remains at zero energy.
@@
Next we study different Fibonacci approximants using the same methodology as above, and find that all approximants with even number of sites follow the same scaling $\sqrt{{\rho_{\mathrm{S}}}}$ in ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$, while the scaling in $\rho$ varies between three different functional forms. Beyond the one shown for the $C_{10}$ approximant, we find ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})=\rho\sqrt{{\rho_{\mathrm{S}}}}$ for e.g. $C_{9}$ in Fig. 19(a), and ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})=\sqrt{{\rho_{\mathrm{S}}}}$ for e.g. $C_{12}$ in Fig. 19(b). The situation changes qualitatively in the approximants with odd number of sites, (i.e. every third approximant $C_{2},C_{5},C_{8},\ldots$), as shown for $C_{11}$ in Fig. 19(c). Here, there is already a zero-energy state in the normal state, related to the microscopic even-odd effect discussed in Sec. III.2, which changes the degeneracy points in the ABS spectrum [151]. As a result, the phase-space of zero-energy ABS ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ technically becomes a manifold instead of an analytic function, where we find zero-energy ABS $\forall{t_{\mathrm{int}}},{\rho_{\mathrm{S}}}$ at $\rho=1$, or alternatively for $\forall\rho,{\rho_{\mathrm{S}}}$ at ${t_{\mathrm{int}}}\to 0$. However, we are not interested in the latter scenario since it describes an uncoupled system. We find that the underlying reason for the broken zero-energy degeneracy, and thus the shift to finite energy, is related to the broken mirror symmetry of the Fibonacci chain, together with the coupling to the superconducting leads (${t_{\mathrm{int}}}\neq 0$). By repeating the above procedure for every approximant from $C_{0}$ to $C_{17}$, we obtain the results presented in Table 1 in Sec. IV.1.

### v2

@@ V.2 Phason angle dependence and topological invariant from critical current
As shown in Fig. 12(c), the subgap states in the Fibonacci chain wind across the gap with the phason angle $\phi$ [17] (see also discussion in Sec. II.2). Our idea to measure the winding is based on that the closer these states wind to the Fermi level, the stronger their contribution is to the current-phase relation, and thus the critical current. Specifically, by tuning the chemical potential close to a gap with winding number $|q|$, the variation with $\phi$ in the spectrum comes predominantly from the winding state, since all other states in comparison show no significant winding (per definition) and therefore have a constant contribution as function of $\phi$. Furthermore, the winding state is at its closest and furthest from the Fermi level exactly $|q|$ times. The conjecture is therefore that this should produce $|q|$ periodic oscillations in the current. We note that this should furthermore even be observable in a small region outside the gap.
Figure 13: (a) Critical current ${I_{\mathrm{c}}}$ as a function of phason angle $\phi$ in a $C_{9}$ junction (no repetition) with ${t_{\mathrm{int}}}={t_{\mathrm{B}}}$, $\rho=0.8$, ${\rho_{\mathrm{S}}}=1$, when the chemical potential $\mu$ lies inside a gap of the Fibonacci chain, and (b) outside the gaps. (c) Power spectrum $P$ of ${I_{\mathrm{c}}}$ (solid) of (a) scaled to show the integer number of periodic cycles $p$ (as $\phi$ varies from 00 to $2\pi$), thus corresponding to the winding number $|q|$ of each gap, $p=|q|$ (dashed). (d) Power spectrum of (b) and Fibonacci numbers $F_{n}$ (dashed). $\mathcal{N}$ is a normalization constant.
@@
To verify the conjecture that the winding numbers $|q|$ appear in the critical current, we plot in Fig. 13(a) ${I_{\mathrm{c}}}$ as a function of $\phi$ at several $\mu$ (line colors) inside the topological gaps, clearly illustrating that ${I_{\mathrm{c}}}$ oscillates periodically $p$ times as $\phi$ varies from 00 to $2\pi$. In contrast, Fig. 13(b) illustrates that when $\mu$ is between the gaps there is no single distinct oscillation frequency [see Fig. 12(c) for gap locations]. To more clearly highlight the periodic oscillations in ${I_{\mathrm{c}}}$, we compute the power spectrum $P$ via the discrete Fourier transform $\mathcal{F}$ with respect to $\phi$
@@
which we normalize with $\mathcal{N}=\sum_{p}P(\mu)$. We also subtract the mean critical current to avoid the trivial peak at $p=0$ in the power spectrum, but refrain from further signal processing operations (e.g. windowing). Figure 13(c) shows the power spectrum as a function of the number of completed periodic cycles $p$ when $\mu$ is inside the gaps (solid), illustrating a single well-defined peak corresponding to the expected winding number, $p=|q|$ (dashed vertical lines). Barely visible are additional peaks mainly at integer multiples of the main peak frequency, e.g. spurious peaks due to the discrete Fourier transform. In contrast, Fig. 13(d) shows that there is no single well-defined peak in the power spectrum when $\mu$ is outside the gaps, but instead a number of small peaks of roughly equal size. Coincidentally, we note that these peaks occur at the Fibonacci numbers $F_{n}\in\{1,2,3,5,8,13,\ldots\}$, with additional smaller peaks at integer multiples of these “Fibonacci peaks” due to the discrete Fourier transform. We consistently find these Fibonacci peaks at other values of $\mu$ and for other junctions (see further below), but have no clear explanation for their appearance, other than that the Fibonacci numbers and golden ratio have a tendency to appear throughout various quantities in the Fibonacci chain [74, 17]. Next, we note that when the Fermi level lies in the middle of the largest gaps ($|q|=1,2,3$), ${I_{\mathrm{c}}}$ vanishes when the state lies at the gap edge, while ${I_{\mathrm{c}}}$ is finite when the winding state is inside the gap, see for instance $\mu=1.57{t_{\mathrm{B}}}$ for the $|q|=3$ gap in Fig. 13(a). These scenarios correspond to when either no state or only the winding state is close to the Fermi level, respectively, thus demonstrating that the topological winding states can fully carry the critical current.
Finally, the above results and conclusions hold more generally for other parameter ranges and approximants, as we demonstrate and discuss in Appendix C.
@@ VI Concluding remarks
We study the influence of quasiperiodicity on the DC Josephson effect, by considering a ballistic hybrid superconductor-quasicrystal-superconductor Josephson junction. We consider the Fibonacci chain as a quasicrystal model system, from the short junction limit with just a few atomic sites, to the long junction limit with several thousand sites. Furthermore, we study Fibonacci chains both without and with repetition, thus essentially modelling both quasicrystals and their approximants [22], embedded across two superconducting leads. As potential experimental realizations of these models, we propose a quasiperiodic engineered atomic chain [118] or metamaterial [161, 162, 163, 164, 165], or moiré structure [121, 122, 123, 124], embedded across two bulk superconductors. Alternatively, we propose a 3D generalization with superconductors sandwiching a Fibonacci superlattice [126, 127, 128, 129, 130], consisting of 2D periodic lattices stacked according to the Fibonacci sequence along the third dimension [126, 127, 128, 129, 130] along which a Josephson current is applied. It was recently shown that superconductivity in the 1D Fibonacci chain extend to such a 3D scenario [73].
@@
We study the Josephson effect via the low-energy ABS spectrum, current-phase relation, and critical current. We exhaustively investigate how these depend on the superconducting phase difference, quasiperiodic degrees of freedom, hopping parameters, chemical potential, and junction length. We find that although the current-phase relation is still $2\pi$ periodic, with either a sinusoidal or sawtooth profile, quasiperiodicity leads to ABS with quasiperiodic oscillations in their probability density, also at perfect resonance. Importantly, we find that quasiperiodicity qualitatively modifies the condition for emergent zero-energy ABS and maximal critical current. We demonstrate that this condition changes between a few simple functional forms in a quasiperiodic manner as the junction length increases. Consequently, we find that the critical current shows quasiperiodic oscillations as a function of junction length, on top of the monotonic decay also found in crystalline junctions [147, 146, 148, 149]. Based on these results we conclude that there might be large sample-to-sample fluctuations between different quasiperiodic junctions, depending on the microscopic details.
@@
Surprisingly, despite proposals for quasiperiodicity enhancing superconductivity and the proximity effect [69, 70, 71, 72, 73, 76, 74, 75], we find that the critical current is not generally enhanced by quasiperiodicity, especially compared to a crystalline junction in the perfect ballistic limit and with a zero-energy state. However, beyond this idealized scenario, we find that quasiperiodicity can significantly enhance the critical current, especially at reduced coupling to the superconducting leads.
@@
We find that by varying the chemical potential in the junction, the junction changes between SNS and SIS behavior in a fractal manner, due to the intrinsic fractal energy spectrum and topological gap structure of the Fibonacci chain [17]. These predictions are directly accessible in experiments, e.g. via an applied gate voltage. Each topological gap hosts a topological subgap state that winds as a function of the phason angle $\phi$, both across the gap and in real space across the Fibonacci chain, with winding number $|q|$ given by a gap labeling theorem [44, 45, 46, 47]. We find that when the chemical potential is tuned through the gaps, these topological subgap states can fully carry the current and that their winding leads to the critical current oscillating with the winding number, while outside the gaps the critical current instead oscillates according to the Fibonacci numbers. We therefore demonstrate how the critical current can in principle measure the topological invariant in the Fibonacci chain. In summary, these results show how Josephson junctions can be used to probe the intricate physics of quasiperiodic systems, including their interplay with ordered states such as superconductivity.
@@
As an outlook, there are many open questions such as the influence of quasiperiodicity on the AC Josephson effect, different heterostructures and sample realizations, the influence of defects and disorder, finite temperature, self-consistent phase gradients, as well as different strengths or symmetries of the superconducting order parameter. It would also be interesting to find a connection with the topological invariant without relying on the phason angle [60]. Another interesting topic is the connection, if any, between the quasiperiodic critical localization with the current and its decay characteristics. Furthermore, other theoretical models like the scattering matrix approach, or analytic calculations using perturbation theory and renormalization group theory [60], might shine additional light on the rich physics of these systems. Beyond quasicrystals, fractal lattices are another example of interesting aperiodic structures to study the interplay of proximitized phase-coherent phenomena [7, 8, 9]. We further note that the supercurrent-magnetic field relation was recently studied in the transverse direction through a single layer of such a fractal [166], i.e. no current was transmitted along fractal degree of freedom itself. Studying the Josephson effect and currents propagating along a fractal would thus also be an interesting outlook.
@@ VII Acknowledgements
Appendix A Microscopic even-odd effect
In this Appendix, we comment on the strong influence of microscopic even-odd effects on the spectrum, specifically in a junction with even or odd number of energy levels, supplementing Secs. IIIV in the main text. The scenario also illustrates what happens in a system where there is already a zero-energy state before the onset of superconductivity. Depending on the actual physical realization of the system, this might either be an idealized or a crucial effect.
In Sec. III we showed how the ABS in the crystalline junction ($\rho=1$) are degenerate at $\Delta\theta=0$ and $2\pi$, and how the degeneracy is lifted for all $\rho\neq 1$ when the quasiperiodic junction breaks the symmetry between the leads. The Fibonacci approximant $C_{9}$ studied in that case has an even number of sites $F_{9}+1=56$, while the number of sites instead becomes odd when repeating the approximant an even number of times, e.g. $1112\times F_{9}+1=$ 111, or in every third approximant e.g. $C_{8}$ with $F_{8}+1=35$ sites. At half-filling, nearest-neighbor hopping endows the chain with chiral symmetry which imposes that for every state with positive energy there exists another state with equal but negative energy. Hence, the normal-state spectrum with an odd number of sites (and thus energy levels) must have an energy level at exactly zero energy due to chiral symmetry, and the ABS spectrum becomes qualitatively modified with the ABS degeneracy points shifting to $\Delta\theta=\pi$ instead [158]. We verify the modified degeneracy in quasiperiodic junctions in Fig. 14, showing the energy-phase spectrum (a) and current-phase relation (b) for a junction consisting of a repeated $C_{9}$ approximant (i.e. otherwise equivalent to Fig. 3). In Fig. 14(a) quasiperiodicity breaks the degeneracy at $\Delta\theta=\pi$ at all higher energies due to the broken symmetry between the leads, but the degeneracy at zero energy is maintained leading to a fully linear phase dispersion with only small variations in slope with $\rho$ at low energy. Consequently, since this zero-energy state carries most of the current, $\rho$ therefore has negligible influence on the current, causing a similar sawtooth profile in all cases shown in Fig. 14(b).
Next, we find that at the critical phase difference $\Delta\theta=\Delta{\theta_{\mathrm{c}}}$, the current is completely carried by the lowest-energy state, where the current contribution from all other states are either exactly zero or pair-wise cancelling. Away from this phase, the reduction in current comes primarily from the variation in the energy-phase dispersion from a linear slope at $\Delta\theta=\pi$ to zero slope at $\Delta\theta=0$ or $2\pi$, and secondarily from a destructive contribution of higher-energy states. This behavior is qualitatively different from the junction with even number of sites described in Sec. III, where the energy-phase dispersion changes also its functional form away from $\Delta{\theta_{\mathrm{c}}}$. Importantly, the zero-energy state in Fig. 14(a) is robust against variations in the hopping parameters. We find that this is related to a preserved degeneracy where e.g. the probability density of the zero-energy state is exactly the same in both leads (not shown), in contrast to the higher-energy states which break the degeneracy.
Appendix B Condition for zero-energy ABS
@@
In this Appendix, we demonstrate how we obtain the condition ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ for the emergence of zero-energy ABS, first in the Fibonacci approximants $C_{n}$ with lengths $L=a_{0}F_{n}$ presented in Table 1, then for more general Fibonacci chains of arbitrary discrete lengths $L$ modeled by the characteristic function Eq. (2), thus supplementing the discussion in Sec. IV.1 and Sec. IV.2. Specifically, we vary simultaneously the hopping parameters ${t_{\mathrm{int}}}$, $\rho\equiv{t_{\mathrm{A}}}/{t_{\mathrm{B}}}$ and ${\rho_{\mathrm{S}}}\equiv{t_{\mathrm{S}}}/{t_{\mathrm{B}}}$ and look for when the lowest-energy ABS reaches zero energy non-trivially (i.e. ignoring trivial cases such as $\rho\to 0$ or ${t_{\mathrm{int}}}\to 0$). We parametrize the phase space with zero-energy ABS as ${t_{\mathrm{int}}}={t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$, where we find a sawtooth current-phase relation and the largest critical current. In a Fibonacci chain with even number of sites, the zero-energy ABS at ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ occurs due to perfect Andreev reflection (zero normal reflection) [146], also with a perfect wave function matching at the interface as evidenced by the probability density of the wave function. In a Fibonacci chain with odd number of sites, there is already a zero-energy state in the normal state [158] (see Appendix A), and ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ corresponds to the parameter space where this state remains at zero energy.
@@
Next we study different Fibonacci approximants using the same methodology as above, and find that all approximants with even number of sites follow the same scaling $\sqrt{{\rho_{\mathrm{S}}}}$ in ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$, while the scaling in $\rho$ varies between three different functional forms. Beyond the one shown for the $C_{10}$ approximant, we find ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})=\rho\sqrt{{\rho_{\mathrm{S}}}}$ for e.g. $C_{9}$ in Fig. 18(a), and ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})=\sqrt{{\rho_{\mathrm{S}}}}$ for e.g. $C_{12}$ in Fig. 18(b). The situation changes qualitatively in the approximants with odd number of sites, (i.e. every third approximant $C_{2},C_{5},C_{8},\ldots$), as shown for $C_{11}$ in Fig. 18(c). Here, there is already a zero-energy state in the normal state, related to the microscopic even-odd effect discussed in Appendix A, which changes the degeneracy points in the ABS spectrum [158]. As a result, the phase-space of zero-energy ABS ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ technically becomes a manifold instead of an analytic function, where we find zero-energy ABS $\forall{t_{\mathrm{int}}},{\rho_{\mathrm{S}}}$ at $\rho=1$, or alternatively for $\forall\rho,{\rho_{\mathrm{S}}}$ at ${t_{\mathrm{int}}}\to 0$. However, we are not interested in the latter scenario since it describes an uncoupled system. We find that the underlying reason for the broken zero-energy degeneracy, and thus the shift to finite energy, is related to the Fibonacci chain breaking the symmetry between the leads, together with the coupling to the superconducting leads (${t_{\mathrm{int}}}\neq 0$). By repeating the above procedure for every approximant from $C_{0}$ to $C_{17}$, we obtain the results presented in Table 1 in Sec. IV.1.